OpenStudy (anonymous):

small leg of a right triangle is 3 inches and hypot is 7 inches. how long is the length of the other leg? give exact and approximation.

OpenStudy (anonymous):

do I use the a^2 + b^2 = c^2 formula

OpenStudy (anonymous):

\[a^2+b^2=c^2\] a=3 inches c=7 inches b=? Plug them in and solve for b. b is the other leg

OpenStudy (anonymous):

yup you are correct.

OpenStudy (anonymous):

remember that c is always the hypotenuse

OpenStudy (anonymous):

what is the difference between the exact and approx?

OpenStudy (anonymous):

so I get b=2√10

OpenStudy (anonymous):

Yes that is the exact answer

OpenStudy (anonymous):

if you try to get the approximate you would plug \[\sqrt{10}\]into a calculator and multiply it to 2

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=2*sqrt%7B10%7D

OpenStudy (anonymous):

so 6.32 it is asking for 2 decimals.

OpenStudy (anonymous):

See where it says decimal approximation?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np your really good at this.

OpenStudy (anonymous):

it has been years......but slowly coming back

OpenStudy (anonymous):

you're*

OpenStudy (anonymous):

yeah man just keep posting if you have more questions

OpenStudy (anonymous):

assume that the vaiable can represent ANY real number √(-4x)^2

OpenStudy (anonymous):

can you help with this one?

OpenStudy (anonymous):

post a new question....I don't really understand it that well.

OpenStudy (anonymous):

see if someone can help you

OpenStudy (anonymous):

ok